OpenStudy (anonymous):

What is the relationship between pi (3,14...) and phi (golden ratio=0.5+sqrt(5)/2=1,618...) ?

OpenStudy (apoorvk):

pi * the golden ratio = 5.08320369

OpenStudy (anonymous):

i am not sure what this question asks. both are irrational, but \[\phi =\frac{1+\sqrt{5}}{2}\] so it is the solution to a quadratic equation, whereas \[\pi\] is not the solution to any polynomial equation (with integer coefficients)

OpenStudy (anonymous):

had to google it Phi = 1 - 2 cos ( 3 Pi / 5) robert everest

OpenStudy (apoorvk):

phi is a historical no.... it sis considered as a golden ratio.. ofcourse other than one of the solutions to a quadratic(that ha sno significance), it is considered the most beautiful no. ever. "1.618" and can be found as a basic for many natural designs, as well in great architectural wonders like the Taj Mahal etc.

OpenStudy (anonymous):

http://www.goldennumber.net/pi-phi-fibonacci.htm